A beam of light is incident at an angle of teta on one face of an equilateral triangular prism made from a material of refractive index 1.6. At what angle of refraction in the prism will the emerging ray graze the opposite face ? What is the time used when the light’s frequency is 4.3 x 10^14 Hz
1
Expert's answer
2019-11-29T09:43:36-0500

In case of minimum deviation, the angle of deviation D is equal to incidence angle i

"D=i (1)"

In our, case

i=60°

Now by Snell’s Law applying at interface of prism:

"\frac { \sin 60°} {\sin r } = 1.6 (2)"


Using (2) we get:

r=33°

The period T is given by formula

"T=\frac {1} {f} (3)"

Using (3) we get:

T=2.3×10-13 s



Answer

33°

2.3×10-13 s